*Find the zero's for the ff quadratic equation:
2x^2-x+8=0




Found 2 solutions by Nate, stanbon:
Answer by Nate(3500)

2x^2 - x + 8 = 0
x^2 - x/2 + 4 = 0
x^2 - x/2 = -4
(x - 1/4)^2 = -4 + 1/16 = -64/16 + 1/16 = -63/16
x - 1/4 = +- i*sqrt(63)/4
x = (1 +- i*sqrt(63))/4

Answer by stanbon(75887)

2x^2-x+8=0
Use the quadratic formula to get:
x=[1+-sqrt(1-4*2*8)]/4
x=[1+-sqrt(-63)]/4
x=[1+-3sqrt(7)i]/4
x=(1+3sqrt(7)i)/4 or x=(1-3sqrt(7)i)/4
-----------------
Cheers,
Stan H.